find the equation of the normal to the parabola y^2+4x=0 at the point where the line y=x+c touches it
1
Expert's answer
2019-03-22T13:55:31-0400

Solution:

First we find the point of intersection of the line


"y=x+c"

with the parabola


"y^2+4x=0:"

"x^2+2cx+c^2+4c=0,"

"x^2+(2c+4)x+c^2=0,"

"D=16(c+1),"

"x_1=-c-2-2\sqrt{c+1},x_2=-c-2+2\sqrt{c+1};""y_1=-2-2\sqrt{c+1},y_2=-2+\sqrt{c+1}"

So


"(x_1;y_1)(,x_2;y_2)"

- wanted points.

Then we have two equations of normal to the parabola at thease points.Find them.

The general equation of normal to the parabola


"y^2=4ax"

at the point t is


"y_0x+2ay-2ay_0-x_0y_0=0"


Then, when

"x_1=-c-2-2\sqrt{c+1}"

and

"y_1=-2-2\sqrt{c+1}"


we hawe the equatin of the normal


"(-2-2\sqrt{c+1})x-2y-6(2+c)-2\sqrt{c+1}(6+c)=0"

or


"y=(-1-\sqrt{c+1})x-3(2+c)-\sqrt{c+1}(6+c),"

when


"x_2=-c-2+2\sqrt{c+1},""y_2=-2+2\sqrt{c+1}"

we hawe the equatin of the normal


"(-2+2\sqrt{c+1})x-2y-6(2+c)+2\sqrt{c+1}(6+c)=0"

or


"y=(-1+\sqrt{c+1})x-3(2+c)+\sqrt{c+1}(6+c)."


Ansver: equations of the normal to the parabola

"y^2+4x=0"

at the point where the line

"y=x+c"

touches it is


"y=(-1-\sqrt{c+1})x-3(2+c)-\sqrt{c+1}(6+c)"

and


"y=(-1+\sqrt{c+1})x-3(2+c)+\sqrt{c+1}(6+c)."